OpenStudy (anonymous):

In an election, 1,860 people voted for candidate A or B. If candidate A received 55% of the votes, how many votes did candidate B receive?

OpenStudy (anonymous):

subtract 55% from 100% which is representing the percentage of the total votes......you get 45%.........that means that candidate B got 45% of the total votes.......

OpenStudy (anonymous):

the choices are (1) 186 (2) 832 (3) 1,023 (4) 1,805

OpenStudy (anonymous):

you could also choose to find 55% of the total votes........and subtract what you get from the total number of votes to find the number of votes candidate B has....

OpenStudy (anonymous):

which method do you prefer??? so I can continue with the explanation..

OpenStudy (anonymous):

the "you could also choose to find 55% of the total votes........and subtract what you get from the total number of votes to find the number of votes candidate B has...."

OpenStudy (anonymous):

I got (3) 1023

OpenStudy (anonymous):

@trini_00x how?

OpenStudy (anonymous):

i can choose to that in just one step..... \[B = 1860- (\frac{ 55 }{ 100 } \times 1860)\] \[B = 837\]

OpenStudy (anonymous):

i think for the the #2 choice it is supposed to be 837 instead of 832......or maybe the value you gave for the total number of votes is wrong...

OpenStudy (anonymous):

Wait, I got 837 votes. I dont know why i got 1023 thats incorrect. But 837 isnt an option...

OpenStudy (anonymous):

Yea I think its supposed to be 837 instead of 832

OpenStudy (anonymous):

yeah sorry! thank you guys so much!

OpenStudy (anonymous):

its 837 not 832! sorry!

OpenStudy (anonymous):

Your welcome :D

OpenStudy (anonymous):

yw...